### v1

# Measurement-induced phase transition in periodically driven free-fermionic systems

Abstract
It is well known that continuous monitoring of a free-fermionic system leads to a steady state entanglement transition from a logarithmically entangled critical to the area law phase. While unitary evolution tends to grow the entanglement, continuous monitoring opposes this growth and results in the pinning of the wavefunction trajectories to the eigenstate of the measurement operators. In this work, we study the fate of this measurement-induced phase transition in a periodically driven free-fermionic quantum system, where the hopping amplitude is varied periodically in time. We show, in the high-frequency limit, that a renormalization group analysis of the non-Hermitian quantum sine-Gordon model [as proposed in Phys. Rev. X 11, 041004 (2021)] reveals a gapless critical phase with logarithmic entanglement entropy growth and a gapped area-law phase, separated by a Berezinskii-Kosterlitz-Thouless (BKT) transition, which is in excellent agreement with our numerical results. Further, the numerical evidence suggests that such BKT transition also prevails in the low-frequency regime, and decreasing frequency tends to favor the critical phase. Strikingly, we also find that if the hopping amplitude is varied completely symmetrically around zero in the form of square-pulse and sinusoidal drive, the BKT transition is absent in the thermodynamic limit, and the system always favors the area-law phase, independent of the frequency regime.
I Introduction
Understanding the entanglement properties of quantum many-body systems has become a key focus of research, particularly in relation to different phases of matter and the transitions between them, as well as identifying quantum chaos at the many-body level Vidmar and Rigol (2017); Vidmar et al. (2017, 2018); Hackl et al. (2019); Murthy and Srednicki (2019); LeBlond et al. (2019); Modak and Nag (2020); Eisert et al. (2010); Modak and Mandal (2021). Thanks to recent progress on the direct measurement of entanglement dynamics with cold atoms and trapped ions, a key question of how entanglement spreads in out-of-equilibrium many-body systems brought lots of attention in the last decade Islam et al. (2015); Kaufman et al. (2016). Entanglement entropy, a quantifier of quantum correlations between subsystems, exhibits distinct scaling behaviors that provide insight into the nature of the underlying phase. For example, in gapped quantum system, the ground state entanglement entropy typically obeys an area law (scales with the boundary area of the subsystem); in contrast, in one-dimensional critical systems, where long-range correlations are prevalent, it follows a logarithmic scaling with subsystem size, reflecting the scale-invariant nature of these systems at criticality Osterloh et al. (2002); Calabrese and Cardy (2004); Pollmann et al. (2009). On the other hand, under unitary evolution, usually for clean short-ranged systems, the entanglement growth is algebraic in time before saturating to a value that scales linearly with the subsystem size, following a volume law Alba and Calabrese (2017); Calabrese and Cardy (2004); Chapman et al. (2019); Modak et al. (2020). However, introducing disorder into systems changes the entanglement behavior drastically. In a closed, non-interacting quantum system, even small amounts of disorder—whether in one dimension or two—can prevent the thermalization of single-particle states, leading to a phenomenon known as Anderson localization Anderson (1958). Furthermore, when interactions are introduced into these disordered systems, the many-body localized (MBL) phase emerges Basko et al. (2006); Oganesyan and Huse (2007); Žnidarič et al. (2008); Pal and Huse (2010); Serbyn et al. (2015); Modak and Mukerjee (2015); Gong et al. (2021); Bardarson et al. (2012a). In the MBL phase, entanglement grows logarithmically over time, despite the absence of energy transport Bardarson et al. (2012b); Serbyn et al. (2013); Prasad and Garg (2022), in stark contrast to thermal systems, where entanglement growth is typically linear Kim and Huse (2013). Moreover, the typical eigenstate entanglement entropy behaves quite differently in the thermal (ergodic) phase, which follows volume law scaling, compared to the non-ergodic MBL phase, where the scaling follows an area law Luitz et al. (2015). Studies of the fate of MBL are not limited to closed systems; there are several interesting results reported where the system is coupled to a bath Wu and Eckardt (2019); Lüschen et al. (2017); Levi et al. (2016); Fischer et al. (2016); Medvedyeva et al. (2016); Hamazaki et al. (2019); Ren et al. (2020).
Another class of entanglement phase transitions arises from the continuous monitoring of quantum systems. The stochastic Schrödinger equation typically governs these open quantum dynamics, and averaging over different realizations of the states leads to the well-known Lindblad master equation dynamics Paule (2018). Since entanglement entropy is a nonlinear function of the state, calculating it for each trajectory separately and then averaging reveals a phase transition. In contrast, any linear function of the state is featureless and does not exhibit such transitions. Several studies have explored these entanglement phase transitions in both interacting and non-interacting systems, considering clean as well as disordered systems with long-range and short-range hopping, using both Lindblad jump operators and quantum state diffusion approaches Alberton et al. (2021); Szyniszewski et al. (2023); Cao et al. (2019); Minato et al. (2022a); Müller et al. (2022); Yamamoto and Hamazaki (2023); Fuji and Ashida (2020); Ladewig et al. (2022). Continuously monitored open systems display competition between unitary evolution, which drives entanglement growth, and measurements, which tend to localize the wavefunction by collapsing it into eigenstates of the measurement operators. As a result, these systems undergo a measurement-induced phase transition as a function of measurement strength, which also has been reported recently in experiments Koh et al. (2023). It has been shown that, in both clean and disordered free-fermionic systems, there is an entanglement phase transition from a critical to an area law phase as the strength of continuous measurements is tuned. This transition falls into the Berezinskii-Kosterlitz-Thouless universality class Cao et al. (2019); Szyniszewski et al. (2023); Alberton et al. (2021), a result that is also supported by a theoretical study in which the dynamics of a continuously monitored free fermionic system were described by an effective non-Hermitian sine-Gordon model Buchhold et al. (2021). There are recent theoretical studies of a monitored chain of a system of size $L$ of Majorana fermions, which shows ($\ln L)^{2}$ behavior in the critical phase Fava et al. (2023). However, it is important to point out that the stability of the critical phase in a one-dimension free-fermionic system under random projective measurements has been questioned recently, and it has been argued in the thermodynamic limit, for any finite measurement strength, only the area-law phase will survive Poboiko et al. (2023). On the other hand, in the case of interacting systems, a transition from a volume law to an area law has been observed Fuji and Ashida (2020); Yamamoto and Hamazaki (2023); Xing et al. (2024). Furthermore, the role of measurements in localizing quantum states in nonunitary quantum circuits has also attracted significant interest Zabalo et al. (2022); Sierant and Turkeshi (2022); Oshima and Fuji (2023); Weinstein et al. (2023); Côté and Kourtis (2022); Block et al. (2022); Sang and Hsieh (2021); Lunt et al. (2021); Li et al. (2021); Fan et al. (2021); Zabalo et al. (2020); Turkeshi et al. (2020); Gullans and Huse (2020a, b); Bao et al. (2020); Choi et al. (2020); Tang and Zhu (2020); Jian et al. (2020). Some of these studies have also been extended in dimension $D>1$  Chahine and Buchhold (2024); Jin and Martin (2024) and a semi-classical limit of such transition has also been introduced recently Ruidas and Banerjee (2024). In contrast to disorder-induced localization observed in the Anderson-localized or MBL phase, the measurement-induced localization is governed by a completely different mechanism and driven by quantum measurements and measurement strength.
In parallel, in the last few decades, periodically driven (Floquet) quantum systems have emerged as a powerful framework for studying non-equilibrium phases of matter Banerjee and Sengupta (2024); Sarkar et al. (2021, 2022); Nag et al. (2014); Agarwala et al. (2016); Luitz et al. (2018). Floquet systems are known to exhibit exotic phases, such as discrete-time crystals Khemani et al. (2016); Else et al. (2016); Zhang et al. (2017) and Floquet topological insulators Oka and Aoki (2009); Lindner et al. (2011); Kitagawa et al. (2011); Kundu et al. (2014); Mukherjee et al. (2016). The fate of many-body localization (MBL) in the presence of a periodic drive has also been studied Lazarides et al. (2015); Ponte et al. (2015). While it has been demonstrated that Floquet systems can stabilize or even induce new phases, it remains an open question how continuous monitoring affects these driven systems and whether the introduction of periodic drive can significantly alter the dynamics of a continuously monitored system, either enhancing or suppressing the effects of measurements. The effect of time-dependent Hamiltonian has been studied within the Lindblad master equation formalism Scopa et al. (2019); Ikeda et al. (2021), but the effect of periodic drive on a continuously monitored open quantum system, particularly in terms of trajectory-averaged entanglement entropy, remains unexplored. In our study, we investigate such cases, and our main finding is that periodically driven non-interacting systems also go through BKT transition. In general, for a given drive protocol, the low-frequency regime favors the critical phase more, implying that the critical measurement strength (transition point) increases with the time period. We also find that the effective dynamics of continuously monitored free-fermionic systems with no drive can mimic the high-frequency limit. Moreover, the symmetric drive of the hopping amplitude about a zero mean is special, such a system always favors the area-law phase in the thermodynamic limit, independent of the frequency regime, in contrast to what has been observed in the case of disorder-induced scenario Sarkar et al. (2021). We summarize our set-up and main results in the form of a schematic in Fig. 1.
This manuscript is structured as follows. In Sec. II, we describe our system and protocols. Then, to build the intuitions about our systems, we focus on a $2\times 2$ toy model in Sec. III. Section. IV is dedicated to the many-body systems, which consists of the analytical prediction for high-frequency limit in sec. IV.1, as well as the numerical results in sec. IV.2. Section. IV.3 concentrates on a particular drive protocol that is symmetric around zero means. Finally, we summarize our results in sec. V.
Figure 1: Schematic of our protocol, where a free-fermionic lattice is continuously monitored, and the hopping amplitude $J(t)$ is periodically varied. Phase diagrams for the critical and area-law phases are shown for different time periods, and $\epsilon/2J=1$ corresponds to no drive result.
II Model and Methodology
We consider spinless fermions in a periodic tight-binding lattice of the length $L$ where the hopping amplitude changes periodically with time, and the system is continuously monitored. The Hamiltonian of this system is given by:
$H(t)=J(t)\sum_{i=1}^{L}\hat{c}_{i}^{{\dagger}}\hat{c}_{i+1}+h.c.,$ (1)
where, $\hat{c}^{{\dagger}}_{i}$ and $\hat{c}_{i}$ are real-space fermionic creation and annihilation operators respectively. We study the effect of two different kinds of driving with the time period (frequency) $T$ ($\omega=1/T$): one is square pulse drive, where $J(t)=+J$ $(-J+\epsilon)$ for $20\leq t<T/$ 2 ($T/2\leq t<T$), with $\epsilon$ varying between 00 to $2J$ (see Fig. 1); the other is the sinusoidal drive where $J(t)=J\sin(2\pi t/T)$. Note that in the limit $\epsilon=0$, the mean of $J(t)$ is 00 over a time period. We examine this limit separately in this manuscript. We primarily present the results for the square pulse in the subsequent sections of the manuscript; only at the end, in Sec. IV.3, some results for the sinusoidal drive are presented. For all our numerical calculations, we choose $J=1/2$. The time evolution of the state can be described using the stochastic Schrodinger equation as
$d|\psi(t)\rangle=-iH(t)dt|\psi(t)\rangle$ $-\frac{\gamma dt}{2}\sum_{i}(\hat{n}_{i}-\langle\hat{n}_{i}\rangle)^{2}|\psi(t)\rangle$
$+\sum_{i}(\hat{n}_{i}-\langle\hat{n}_{i}\rangle)dW_{i}^{t}|\psi(t)\rangle.$
The above equation describes the continuous monitoring of the particle number $\hat{n}_{i}=\hat{c}_{i}^{{\dagger}}\hat{c}_{i}$ on each site, with measurement strength $\gamma$ Cao et al. (2019). $dW_{i}$s are taken from uncorrelated normal distributions having mean 00 and variance $\gamma dt$, which implies, $\overline{dW_{i,t}}=0$     and    $\overline{dW_{i,t}dW_{j,t^{\prime}}}=\gamma dt\delta_{i,j}\delta(t-t^{\prime})$. For our model and protocols, the total number of particles $N=\sum_{i}\langle\hat{n}_{i}\rangle$ remain conserved in the entire time evolution process; hence, if we start with an initial state with a particular filling at $t=0$, for any time $t>0$, the filling fraction remains the same. In our work, we have fixed the filling fraction $N/L=1/2$ for all many-body calculations. Also, note that the last two terms in Eq. (LABEL:nlse) describe the contribution from the continuous measurement process and can be realized by homodyne detection in quantum optics Wiseman and Milburn (1993); Øksendal and Øksendal (2003).
In order to solve Eq. (LABEL:nlse), one needs to perform Trotterization and can write,
$|\psi(t+dt)\rangle=\mathcal{N}e^{\mathcal{M}}e^{-iH(t)dt}|\psi(t)\rangle,$ (3)
where, $\mathcal{M}=\sum_{i}[dW_{i}^{t}+(2\langle\hat{n}_{i}\rangle-1)\gamma dt]\hat{n}_{i}$, and $\mathcal{N}$ is the normalization constant. Any pure Gaussian state can be represented as,
$|\psi\rangle=\prod_{i=1}^{N}(\sum_{j=1}^{L}U_{ji}\hat{c}_{j}^{\dagger})|0\rangle.$ (4)
As $|\psi(t)\rangle$ evolves according to Eq. (3), and it preserves the Gaussianity of the state, hence; it is straight forward to show that the $U$ evolves as,
$U(t+dt)=e^{M}e^{-iH(t)dt}U(t),$ (5)
where $H(t)=J(t)\sum_{i=1}^{L}(\hat{c}_{i}^{{\dagger}}\hat{c}_{i+1}+h.c.)$ is the single particle Hamiltonian and $M_{ij}=\delta_{i,j}[dW_{i}^{t}+(2\langle\hat{n}_{i}\rangle-1)\gamma dt]$. Under this non-Hermitian evolution, we need to properly normalize the wavefunction in each step, which can be done using the $QR$ decomposition of the $U$ matrix in each step as $U(t+dt)=QR$, and we set the new $U$ as $Q$. Note that $U$ is an $L\crossproduct N$ matrix, follows $U^{{\dagger}}U=\mathbb{I}_{N\crossproduct N}$. On the other hand, the correlation matrix $D$ can be defined as $D_{ij}=\langle\hat{c}_{i}^{{\dagger}}\hat{c}_{j}\rangle=[UU^{{\dagger}}]_{ij}$ For our many-body calculations, we choose $dW_{i}$s from different random realizations, compute the nonlinear function of states, e.g., entanglement entropy using the correlation matrix for each realization separately, and finally average over different realizations (we ensure the convergence of our results by varying the realization number). On the other hand, the evolution of the average state will be governed by the Lindblad equation Cao et al. (2019).
III $2\times 2$ Toy model
Figure 2: The variation of the expectation value of fermionic occupancy, $\langle\hat{n}_{1}\rangle$, at site 1 as a function of time $t$ for a typical single trajectory in a $2\times 2$ periodically driven toy model, for different values of $T$ and with $\epsilon=0$ fixed.
Figure 3: $\overline{S_{E}(\gamma)}$ vs.$\gamma$ plots for different driving periods for $\epsilon=0$ at steady state. The black dashed line corresponds to no-drive results with $J(t)=J=1/2$.
Figure 4: $\overline{S_{E}(\gamma)}$ vs.$\gamma$ plots for different $\epsilon$ values, for $T=0.5$ at steady state. In the limit $\epsilon=2J$, we get back the no-drive($J(t)=J=1/2$) result, which is represented by the black dashed line.
In this section, we study a periodically driven, continuously monitored $2\times 2$ system to gain insight into the effect of driving on the system under a square pulse driving protocol. We use the same Hamiltonian as in Eq. 1 and the same protocol as described in Sec. II but limit ourselves to two sites and one particle. In this system, the Hilbert space is spanned by the basis states $|01\rangle$ and $|10\rangle$, which are also the eigenstates of the measurement operators $\hat{n}_{1}$ and $\hat{n}_{2}$. In Fig. 2, we plot the expectation value of the occupancy $\langle\hat{n}_{1}\rangle$ at site $1$ for a typical single trajectory as a function of time, for different driving periods $T$, but for the fixed $\gamma=0.1$ and $\epsilon=0$. For very small $T$, i.e., in the high-frequency regime, $\langle\hat{n}_{1}\rangle$ remains mostly at either 00 or $1$, indicating that the particle is pinned or localized in the eigenstates of the measurement operator or at the dark states, even for a small measurement strength $\gamma=0.1$. In contrast, as we increase $T$, the particle becomes depinned from the dark states.
To quantify this behavior more systematically, we define a measure,
$S_{E}(\gamma)={-\langle\hat{n}_{1}\rangle\ln\langle\hat{n}_{1}\rangle-\langle\hat{n}_{2}\rangle\ln\langle\hat{n}_{2}\rangle}.$ (6)
$S_{E}(\gamma)$ is averaged over different random trajectories, followed by a long-time average, which we denote as $\overline{S_{E}(\gamma)}$. In the localized or pinned phase, where both $\langle\hat{n}_{1}\rangle$ and $\langle\hat{n}_{2}\rangle$ remain at 00 or $1$ most of the time, the steady-state value of $\overline{S_{E}(\gamma)}$ tends to zero. In contrast, in the delocalized phase, this value becomes significantly greater than zero. In Fig. 3, we plot the steady-state value of $\overline{S_{E}(\gamma)}$ as a function of $\gamma$ for different driving periods. For a very small driving period, $T=1$, $\overline{S_{E}(\gamma)}$ is almost zero for all $\gamma$. As we increase $T$, the steady-state value $\overline{S_{E}(\gamma)}$ gradually increases at-least for small $\gamma$. This behavior indicates a potential transition from depinning to pinning as $\gamma$ increases, particularly for large $T$. Interestingly, for $T\gtrsim 2.5$, the steady-state value $\overline{S_{E}(\gamma)}$ stabilizes and shows almost negligible change upon increasing $T$ further. Moreover, $T\gtrsim 2.5$ data is almost indistinguishable from the no-drive results. This gives us a hint that large $T$ promotes depinning, and also, for $\epsilon=0$, the large $T$ limit may resemble a no-drive scenario even for many-body systems.
While Fig. 3 focuses on the case with $\epsilon=0$, we now examine the dependence of $\overline{S_{E}(\gamma)}$ on $\epsilon$ in the high-frequency regime with $T=0.5$. Figure 4 shows that for $\epsilon=0$, $\overline{S_{E}(\gamma)}$ remains nearly zero for all $\gamma$, as also observed in Fig. 3. As $\epsilon$ increases, we observe a typical increase in $\overline{S_{E}(\gamma)}$. However, with further increases in $\gamma$, $\overline{S_{E}(\gamma)}$ approaches zero even for finite $\epsilon$, indicating a transition from depinning to pinning. From Fig. 4, it is also evident that the decay of $\overline{S_{E}(\gamma)}$ to zero occurs more rapidly for smaller $\epsilon$ compared to larger $\epsilon$. This suggests, indirectly, that the critical measurement strength $\gamma_{c}$ (if a measurement-induced transition exists) may decrease as $\epsilon$ decreases, favoring more the localized phase. In the following sections, we will now focus on the many-body Hamiltonian.
IV Many-body system
In this section, we focus on the periodically driven many-body system described by the Hamiltonian in Eq. (1), which is continuously monitored. We initialize the system in a Néel state $|\psi_{0}\rangle=|1010\cdots\cdots 1010\rangle$ at time $t=0$ and let the system evolve according to Eq. (LABEL:nlse). In the absence of a drive, such a system undergoes a transition from the critical to the localized phase (depinning to pinning transition), which has been previously characterized using the scaling of entanglement entropy with system size Skinner et al. (2019); Minato et al. (2022b); Fava et al. (2023). In our work, we also use entanglement entropy as a diagnostic tool.
The entanglement between subsystem $A$ of length $l_{A}$ and the rest of the system is measured by the von Neumann entanglement entropy, which is defined as $S=-\text{tr}(\rho_{A}\ln\rho_{A})$ Amico et al. (2008), where the reduced density matrix $\rho_{A}$ is obtained by tracing over the degrees of freedom of the rest of the system (which we identify as $B$), i.e., $\rho_{A}=\text{tr}_{B}|\psi(t)\rangle\langle\psi(t)|$. Even in the continuously monitored driven system we are interested in, the evolution generators are quadratic, preserving the Gaussianity of the quantum trajectories. Hence, the entanglement entropy $S$ can be easily calculated using the eigenvalues $\lambda_{i}$ of the correlation matrix $D_{ij}=\langle\hat{c}_{i}^{{\dagger}}\hat{c}_{j}\rangle$ as Peschel (2003); Peschel and Eisler (2009); Vidmar et al. (2017),
$S(l_{A},L)=-\sum_{i=1}^{l_{A}}\left[\lambda_{i}\ln\lambda_{i}+(1-\lambda_{i})\ln(1-\lambda_{i})\right].$
In our calculation, we take $l_{A}=L/2$, where $L$ is the total system size. Thus, the eigenvalues of the correlation matrix $D_{ij}$ are computed by restricting the indices up to $L/2$. After calculating the entanglement entropy for each trajectory, we average over different realizations, which we denote as $\overline{S}$. In the subsequent sections, we study the behavior of $\overline{S}$ in great detail for different frequency regimes.
Figure 5: Flow of $\lambda$ for different $\epsilon$ values, with same initial condition $|\lambda=0.1|$ and $|K(\epsilon=2J)|=1.42\pi$. For $\epsilon=2J$, the flow of $\lambda$ is towards the Gaussian fixed point, which suggests a critical phase; on decreasing $\epsilon$, the flow of $\lambda$ approaches the strong coupling fixed point, which suggests area-law entanglement phase. In our calculation, we take $A=1$.
IV.1 High-frequency regime: Analytical prediction
It has been proposed recently that in the absence of the driving, the long-time dynamics of the continuously monitored fermionic Hamiltonian given by Eq. (1) could be theoretically described by an effective bosonized non-Hermitian sine-Gordon model  Buchhold et al. (2021),
$H_{eff}=\frac{vJ}{2\pi}\int_{x}\bigg{[}(\partial_{x}\hat{\theta}_{x})^{2}+(1-\frac{2i\gamma}{vJ\pi})(\partial_{x}\hat{\phi}_{x})^{2}\bigg{]}$
$+i\lambda\int_{x}\bigg{[}\cos(\sqrt{8}\hat{\phi}_{x}-1)\bigg{]}.$ (7)
Here, $v=\partial_{k}E(k)|_{k=k_{F}},E(k)=2\cos k$ for free theory and $k_{F}$ is fermi momentum, $\hat{\phi}_{x}$ is the bosonic field operator with its conjugate field $\hat{\theta}_{x}$ obeying the commutation relation $[\partial_{x}{\hat{\theta}_{x}},\hat{\phi}_{x^{\prime}}]=-i\pi\delta(x-x^{\prime})$, and $\lambda=\gamma m^{2}$, where $m$ is $O(1)$, depends on normal ordering prescription. It has been shown recently that in a long-time steady state, the system comes to a dark state, which is defined as the zero energy eigenstate of the $H_{eff}$, i.e., $H_{eff}|\psi_{D}\rangle=0$ (or equivalently can be identified as an eigenstate of the evolution operator $U(T)=e^{-iH_{eff}T}$ with eigenvalues $1$, i.e., $U(T)|\psi_{D}\rangle=|\psi_{D}\rangle$). The steady-state entanglement entropy can be obtained from  Casini and Huerta (2009); Calabrese and Cardy (2004),
$S(l_{A})=\frac{1}{3}\langle\psi_{D}|\hat{\phi}_{x}\hat{\phi}_{x+l_{A}}|\psi_{D}\rangle.$ (8)
In the case of the square pulse driving protocol (as discussed in Sec. II), in a similar spirit, one can write down the effective Floquet evolution operator in terms of the non-Hermitian sine-Gordon Hamiltonian as
$U_{F}(T)=e^{-iH_{eff}^{F}T}=e^{-iH_{eff}^{-}T/2}e^{-iH_{eff}^{+}T/2},$ (9)
where,
$H_{eff}^{+}=\frac{Jv}{2\pi}\int_{x}\bigg{[}(\partial_{x}\hat{\theta}_{x})^{2}+(1-\frac{2i\gamma}{Jv\pi})(\partial_{x}\hat{\phi}_{x})^{2}\bigg{]}$
$+i\lambda\int_{x}\bigg{[}\cos(\sqrt{8}\hat{\phi}_{x}-1)\bigg{]}.$
and
$H_{eff}^{-}=\frac{(-J+\epsilon)v}{2\pi}\int_{x}\bigg{[}(\partial_{x}\hat{\theta}_{x})^{2}+$ $(1-\frac{2i\gamma}{(-J+\epsilon)v\pi})(\partial_{x}\hat{\phi}_{x})^{2}\bigg{]}$
$+i\lambda\int_{x}\bigg{[}\cos(\sqrt{8}\hat{\phi}_{x}-1)\bigg{]},$
and in the high-frequency limit, the effective Floquet Hamiltonian $H_{eff}^{F}$ can be calculated using Magnus expansion as Bukov et al. (2015),
$H_{eff}^{F}=\frac{1}{2}(H_{eff}^{-}+H_{eff}^{+})-\frac{iT}{4}[H_{eff}^{-},H_{eff}^{+}]+O(T^{2}).$ (10)
In the extremely high-frequency limit $(T<<1)$, $H_{eff}^{F}$ can be approximated by the 1st term in the Magnus expansion as,
$H_{eff}^{F}\simeq\bigg{[}$ $\frac{v\epsilon}{4\pi}\int_{x}\bigg{[}(\partial_{x}\hat{\theta}_{x})^{2}+\eta^{2}(\partial_{x}\hat{\phi}_{x})^{2}\bigg{]}$
$+i\lambda\int_{x}\bigg{[}\cos(\sqrt{8}\hat{\phi}_{x}-1)\bigg{]},$ (11)
where $\eta^{2}=1-\frac{4i\gamma}{\epsilon\pi v}$ Hence, one would expect that in such a high-frequency limit, the steady state of the driven continuously monitored system, as discussed in Sec. II, can be described by the dark state of the non-Hermitian sine-Gordon Hamiltonian Eq. (11) and steady state entanglement entropy can be obtained using Eq. (8), $|\psi_{D}\rangle$ correspond to the zero energy eigenstate of the Hamiltonian Eq. (11). Moreover, this also suggests that the entanglement dynamics in the low-frequency limit can effectively be mimicked by the continuously monitored no-drive dynamics of tight-binding Hamiltonian with effective hopping amplitude $J_{eff}=\epsilon/2$. It also automatically implies if in the no drive case, i.e., $\epsilon=2J$, the transition point is at $\gamma_{c}=\gamma_{0}$, for $\epsilon<2J$ (the drive is introduced), the transition point will be $\gamma_{c}=\epsilon\gamma_{0}/2J$, decays to zero linearly with decreasing $\epsilon$.
In the case of the $\epsilon>0$ and $\lambda=0$, the Hamiltonian Eq. (11) has a Gaussian fixed point, which implies the logarithmic scaling of the entanglement entropy with sub-system size will prevail Buchhold et al. (2021), i.e.,
$S=\frac{1}{3}C({\gamma})\ln(l_{A}),$ (12)
where $C({\gamma})$ is referred to as the effective central charge. We refer to this phase as a critical phase. One needs to find out renormalization flow equations to study the effect of the nonlinearity $\lambda$ and investigate the fate of the Gaussian fixed point. The Hamiltonian Eq. (11) can be brought in its action form using complex wick rotation $(x,t)\to({\tilde{\epsilon}}^{1/2}\tilde{\eta}^{1/2}x,i\tilde{\epsilon}^{-1/2}\tilde{\eta}^{-1/2}t)$, and the action $\mathcal{F}$ reads as,
$\mathcal{F}=\int_{X}\frac{K}{16\pi}(\gradient\hat{\phi}_{X})^{2}+i{\lambda}\cos(\hat{\phi}_{X}),$ (13)
where $(\gradient\phi_{x})^{2}=(\partial_{x}\hat{\phi}_{X})^{2}+(\partial_{t}\hat{\phi}_{X})^{2},X\in(x,t)$, $\tilde{\epsilon}=v\epsilon/2\pi$, and $K^{2}=\eta^{2}$. We use a similar renormalization group (RG) prescription as discussed in Ref. Buchhold et al. (2021); Amit et al. (1980), first, decompose the fields into short-range modes that correspond to momentum $\frac{\Lambda}{\zeta}<|k|<\Lambda$ and the long-range modes that correspond to $|k|<\frac{\Lambda}{\zeta}$ ($\Lambda$ is a short-distance cutoff, and $\zeta=e^{s}$, and $s$ is the rescaling parameter controlling the renormalization group flow), and finally, integrate out the short-range modes. One can write down renormalization group flow equations perturbatively in ${\lambda}$ as (note these equations are identical as obtained in Ref. Buchhold et al. (2021)) ,
$\partial_{s}\lambda=(2-\frac{8\pi}{K})\lambda$
$\partial_{s}K=-\lambda^{2}A,$ (14)
where $A$ is a positive number of order $O(1)$ determined by the propagator of the Gaussian theory. If increasing the RG iteration step $\lambda$ grows, it signifies that the non-linearity is relevant perturbation; on the other hand, if $\lambda$ decays, that will make the perturbation irrelevant, and the physics will still be governed by the Gaussian fixed point, implies the survival of the critical phase. In Fig. 5 we numerically solve the Eq. (14) with initial condition $|\lambda(s=0)|=0.1$ and $K(s=0,\epsilon)=\sqrt{\frac{2J}{\epsilon}{K^{2}(s=0,\epsilon=2J)}+1-\frac{2J}{\epsilon}}$, with $|K(s=0,\epsilon=2J)|=1.42\pi$. We choose the initial condition in such a way that for $\epsilon=2J$ (no-drive scenario), $\lambda$ decreases with increasing $s$, signifies RG flow towards the Gaussian fixed point, implying the critical phase. The question we ask is if we change $\epsilon$ (indirect way to introduce drive), how $\lambda$ will change with RG steps? We find that for $\epsilon=0.75J$ and $0.55J$, it decreases with increasing $s$. On the other hand, on decreasing $\epsilon$ further, the $\lambda$ starts growing with $s$, which makes the non-linear $\lambda$ term relevant. It signifies that physics is no longer governed by the Gaussian fixed point; instead, it is governed by a strong coupling fixed point, which is responsible for the area-law entanglement phase. This study also shows that decreasing $\epsilon$ promotes the area-law entanglement phase, and if one starts with an initial condition such that the no-drive scenario shows a critical phase, with decreasing $\epsilon$ one should be able to probe the critical-area law phase transition Alberton et al. (2021).
Next, we try to understand the effect of $\epsilon$ on the critical phase, more precisely, how the correlation or entanglement entropy depends on our drive protocol. We just focus on the $\lambda=0$ limit of the Hamiltonian  (11), which can be expressed in the momentum space representation as,
$H^{F}_{eff}(\lambda=0)=\frac{v\epsilon}{4\pi}\int_{q}q^{2}\bigg{[}\hat{\theta_{q}}\hat{\theta}_{-q}+\tilde{\eta}^{2}\hat{\phi}_{q}\hat{\phi}_{-q}\bigg{]}.$
It has been argued that this Hamiltonian has a unique dark state Buchhold et al. (2021), and the correlation function in this dark state is given by,
$\langle\psi_{D}|\hat{\phi}_{q}\hat{\phi}_{-q}|\psi_{D}\rangle=\frac{v^{3/2}\sqrt{\pi\epsilon}}{8\gamma|q|}\sqrt{2\sqrt{16\gamma^{2}+{(v\pi\epsilon)}^{2}}-2v\pi\epsilon}.$
It is straightforward to check that for a fixed non-zero $\gamma$, in the limit $\epsilon\to 0$, this correlation tends to zero, and the effective central charge $C(\gamma)$ (note that the leading order term in the entanglement entropy $S=\frac{C(\gamma)}{3}\ln l_{A}$) will also approach zero as $\sqrt{\frac{\epsilon}{\gamma}}$. This implies that in the case of $\epsilon=0$, for any finite $\gamma$, one would expect to observe the area-law phase. In a nutshell, we make three main predictions about the high-frequency limit drive: 1) the entanglement dynamics can be mimicked by the dynamics of no-drive continuously monitored free-fermionic chain with the hopping amplitude $J_{eff}=\epsilon/2$, 2) If one starts in critical-phase in the no-drive regime, i.e. $\epsilon=2J$, with decreasing $\epsilon$, one can be able to probe critical to area-law phase transition, 3) $\epsilon=0$ limit will not display entanglement transition, it will always show the area-law entanglement for any non-zero finite measurement strength, 4) the transition point $\gamma_{c}$ is expected to decay to zero linearly with decreasing $\epsilon$. Next, we will try to test our analytical prediction using numerical simulation on a finite-size tight-binding lattice.
Figure 6: Stroboscopic growth of trajectory averaged entanglement entropy for $\gamma=0.05$ (upper panel) and $\gamma=0.20$ (lower panel). All plots are for $\epsilon=3J/4$ and $T=5$. The upper panel inset shows that in the high-frequency regime $T=0.5$, the numerical results are in agreement with no-drive results for $J_{eff}=3J/8$. Inset data is for $\gamma=0.05$ and $N=250$.
Figure 7: Upper Panel: System size scaling of steady state trajectory averaged entanglement entropy $\overline{S(L)}$ for different $\gamma$. Lower Panel: Data collapse of entanglement entropy for different $L$ and $\gamma$ assuming the BKT type scaling ansatz with critical measurement strength $\gamma_{c}=0.086$. The results are for $T=5$ and $\epsilon=3J/4$. The inset in the lower panel shows the variation of the $\gamma_{c}$ (obtained from the data-collapse) for different $\epsilon$.
Figure 8: Color variation corresponds to the different values of the effective central charge $C(\gamma)$ for $T=5$. Black Dots represent $\gamma_{c}$, and the dashed line is the best-fit line, identifying the phase transition.
IV.2 Numerical results
In this section, we perform a numerical experiment on a finite-size lattice with periodic boundary conditions using the methodology described in Sec. II. First, we study the time evolution of the entanglement entropy, i.e., averaged over many different trajectories, starting from the initial zero entanglement product state for different values of $\epsilon$ and measurement strengths $\gamma,$ for $T=5$. Figure. 6 shows the dynamics of $\overline{S(nT)}$ for $\epsilon=3J/4$. The upper panel shows for $\gamma=0.05$, the saturation value of the entanglement entropy increases with system size; on the other hand, the lower panel shows for $\gamma=0.2$, the steady state value of the average entanglement entropy hardly changes with system size. This is a signature that, with increasing measurement strength, an entanglement transition may occur, where for small values of $\gamma$, the steady-state entanglement increases with $L$, and for large $\gamma$, it does not scale with $L$. The inset of the upper panel shows the entanglement dynamics in high-frequency regime $T=0.5$ for $\gamma=0.05$ and $L=250$ (red-dotted line). As predicted in the previous section, we find that, indeed, for $T=0.5$, the entanglement dynamics can be mimicked by the continuously monitored no-drive scenario in a tight-binding lattice with an effective hopping strength $J_{eff}=\epsilon/2$, which is denoted by points in the inset figure. We further check the system size scaling of steady-state entanglement entropy for $\epsilon=3J/4$ in Fig. 7 (upper panel) for different $\gamma$ values. It shows clear evidence of a logarithmic-area law entanglement phase transition. To re-confirm that for small $\gamma$, the scaling is logarithmic (not volume law), we also perform an F-test (see Appendix. C). It has been shown recently that in the absence of driving, the measurement-induced critical to area-law phase transition with increasing $\gamma$ belongs to the BKT university class Szyniszewski et al. (2023); Alberton et al. (2021). The system size dependence of the entanglement entropy in the BKT universality class obeys the following scaling form in the vicinity of the critical $\gamma_{c}$ Harada and Kawashima (1997),
$\bar{S}(L/2,L,\gamma)-\bar{S}(L/2,L,\gamma_{c})=F[(\gamma-\gamma_{c})(\ln L)^{2}].$ (15)
Given that the steady entanglement entropy growth is logarithmic in a small $\gamma$ regime, we also use the similar BKT scaling ansatz to obtain the critical $\gamma_{c}$ values for different $\epsilon$ and different values of $T$. Figure. 7 (lower panel) shows the data collapse of steady state entanglement entropy for different $L$ and $\gamma$, which confirms that the transition belongs to the BKT universality class and allows us to evaluate the critical $\gamma_{c}$. We use the cost function minimization technique to obtain the $\gamma_{c}$ (see Appendix. B). Inset shows the critical $\gamma_{c}$ values for different $\epsilon$. Our results suggest the critical $\gamma_{c}$ decreases with decreasing $\epsilon$ and approaches zero in the $\epsilon\to 0$ limit, as also predicted by our previous RG study for high-frequency limit. The best fit suggests $\gamma_{c}\sim(\epsilon/2J)^{\nu}$, were $\nu\simeq 0.91$, slightly smaller than $1$. Note that $\nu=1$ was predicted previously in sec. IV.1 for the high-frequency limit. Figure. 8 shows the contour plot of the effective central charge $C(\gamma)$ for different values of $\gamma$ and $\epsilon$, and dots correspond to the $\gamma_{c}$ and the dashed line corresponds to the best-fit line. Also, important to point out, given our $T=0.5$ data are almost indistinguishable from the no-drive data with $J_{eff}=\epsilon/2$, one will obtain $\nu=1$ for $T=0.5$. $\nu<1$ for $T=5$ also signifies that with increasing $T$, $\gamma_{c}$ will increase for a given $\epsilon$; it will promote the critical phase more. Moreover, the data for $T=5$ also suggests that for a symmetric drive with zero means, i.e., for $\epsilon=0$, $\gamma_{c}\to 0$, implies no transition, always area-law phase. Though our analytical results suggest the same for the high-frequency regime; it is not at all obvious that this phenomenon persists even for large $T$. Especially given that there are examples where disorder-induced localization-delocalization transition has been observed in the large $T$ limit for symmetric drive mean around zero Sarkar et al. (2021, 2022). Hence, in the next section, we investigate the $\epsilon=0$ limit even more carefully.
Figure 9: Stroboscopic growth of trajectory averaged entanglement entropy for different $T$ and $\gamma$. Upper panel: for $T=20$ and varying $\gamma$, Lower panel: for $\gamma=0.05$ and varying $T$. Plots are for system size $L=250$ and $\epsilon=0$.
Figure 10: Upper panel: Scaling of the steady state entanglement entropy with system size $L$ for different $T$, but fixed $\gamma=0.05$ and $\epsilon=0$. Initially, it shows logarithmic growth with $L$ and then tends to saturate. Dashed-line corresponds to no-drive $J=1/2$ results for $\gamma=0.05$. Lower panel: plot of $L_{c}$ vs. $T$, $L_{c}$ is estimated using Eq. (16), circles are for square pulse protocol, and triangles are for sinusoidal protocol. Inset shows the best fit of the entanglement entropy assuming the functional form given by Eq. (16) for square pulse protocol.
IV.3 Symmetric drive with zero mean: $\epsilon=0$ limit
Square pulse drive: Here, we focus on the $\epsilon=0$ limit. First, the time evolution of the entanglement entropy, i.e., averaged over many different trajectories, starting from the initial Neel state for different values of time period $T$ and measurement strengths $\gamma,$ is studied. We find that $\overline{S}$ increases with time initially and then saturates to a steady state value. From Fig. 9 upper panel, we see, for a fixed driving period $T=20$, if we increase the measurement strength $\gamma$, the steady state entanglement saturation value is decreased; similar to our previous results, the entanglement spread gets suppressed because of increasing measurement strength. On the other hand, it is shown in the lower panel of Fig. 9 that, for a fixed measurement strength $\gamma=0.05$, the entanglement saturation value increases with the increase of the driving time-period $T$. This suggests, in general, that the low driving frequency promotes the depinning of the particles, and the entanglement entropy saturation values approach the no-drive result with increasing $T$. We also noticed there is a critical driving period $T_{c}$, such that for $T>T_{c}$, the results are almost indistinguishable from those of the no-drive scenario. However, this $T_{c}$ typically increases with $L$. The fact that $T>T_{c}$ resembles a no-drive situation has also been observed for our $2\times 2$ Toy model previously.
Next, we immediately ask the following question: How does the steady-state entanglement entropy change with the subsystem size for different $T$? The existence of a $T_{c}$ for each $L$ might be a tempting scenario to conjecture that in the very large $T$ limit, $\epsilon=0$ results should resemble no-drive results and should display an entanglement transition. However, one needs to take the large $T$ and large $L$ limits carefully. The order of these two limits doesn’t commute. If one takes a large $T$ limit first, then any finite $L$ (no matter how large it is), $T_{c}$ has been achieved, and hence, one will observe no-drive results, i.e., the presence of the measurement-induced transition. On the other hand, the correct limit is to take the large $L$ limit first, then increase the $T$. In that case, the outcome will be different, and we further try to argue that in this limit, there will be no transition.
Given for each $L$, a critical $T_{c}$ exists, and if one chooses a reasonably large $T$, it will turn out to be a $T_{c}$ for some system size $L=L_{1}$. Hence, for that particular $T$, till $L<L_{1}$, the saturation value will resemble the no-drive results, which implies it will show logarithmic scaling for $L<L_{1}$ for small $\gamma$, but for $L>L_{1}$, the steady entanglement data vs $L$ will certainly show a bending because if $T<T_{c}$, the entanglement saturation value is smaller than the $T\geq T_{c}$ value. We precisely see that in Fig. 10 upper panel. We find for $\gamma=0.05$ and $T=5$, and $8$, the steady state entanglement entropy is almost constant for all $L$, which suggests an area law scaling with the system size. On the other hand, as we increase $T$, we see an initial logarithmic scaling of entanglement entropy with system size till $L<L_{1}$, followed by bending trend, finally after a critical $L=L_{c}$, entanglement starts showing area law behavior. The dashed line corresponds to the no-drive results for $J=1/2$ and $\gamma=0.05$. Roughly, till $L\simeq 20$, this no-drive results resembles $T=20$ data, beyond $L=20$, the no-drive results keep increasing with $L$, while the $T=20$ data starts bending (see appendix. A for $T=100$ results). To estimate the $L_{c}$, we fit the steady entanglement entropy vs $L$ data with a two parameters ($S_{sat}$ and $L_{c}$) fitting function $f(L)$, i.e.,
$f(L)=S_{sat}\tanh\bigg(\frac{\ln L}{\ln L_{c}}\bigg{missing}),$ (16)
in the inset of Fig. 10 lower panel. In the limit $L>>L_{c}$, $f(L)\to S_{sat}$, which is a constant suggesting area-law phase, on the other hand, for $L<<L_{c}$, $f(L)\propto\log L$. We estimate $L_{c}$ from the best fit and plot $L_{c}$ vs $T$ in Fig. 10 lower panel, which suggests an exponential growth of $L_{c}$ with the driving period $T$. As an exercise, we estimate the fitting parameter $L_{c}$ for $T=100$ and $\gamma=0.05$ from the exponential fit, which gives $L_{c}\simeq 10^{16}$. Hence, it is almost impossible to reach the accessible system size regime to see area law scaling for $T=100$ (see Fig. 12 in Appendix A for details).
Sinusoidal drive: To become more certain about our findings for the $\epsilon=0$ limit, we also study the effect of sinusoidal driving, where $J(t)=J\sin(2\pi t/T)$. We find that even for the sinusoidal drive protocol, the area-law entanglement phase sustains in the large $L$ limit. Growth of trajectory averaged entanglement entropy for $T=20$ and $\gamma=0.05$ are shown in Fig. 11 for different $L$. We find that $L=400$ and $550$ data are almost indistinguishable, though for small $L$, the saturation values are different. The inset of Fig. 11 shows the variation of the steady state entanglement entropy with $L$ for different $T$, and it clearly shows the evidence of the area law in the large $L$ limit, as already observed for the square pulse case. We also estimate the $L_{c}$ using the fitting function $f(L)$ (see Eq. (16)) and compare it with square pulse results in Fig. 10 lower panel. We find that $L_{c}$ is even smaller for sinusoidal drive compared to the square pulse.
Figure 11: Stroboscopical growth of trajectory averaged entanglement entropy for different $L$ for the sinusoidal driving protocol for $T=20$ and $\gamma=0.05$(main panel). Inset shows the system size scaling of the steady state trajectory averaged entanglement entropy for the different driving periods $T$. Data presented here are for $\gamma=0.05$.
V Conclusions
In this work, we investigate the effect of periodic drive in a continuously monitored free-fermionic lattice model. It is well-known that in the absence of the drive, such systems undergo a measurement-induced entanglement phase transition from critical (logarithmic) to area law. Our main finding is that periodically driven non-interacting systems also go through a measurement-induced entanglement phase transition, and the universality class of this transition also seems to be preserved, which belongs to the BKT universality class. In general, for a given drive protocol, the low-frequency regime favors the critical phase more, implying that the critical measurement strength (transition point) increases with the time period. Moreover, we have also investigated a drive protocol in which the hopping amplitudes are varied completely symmetrically around zero in the form of square-pulse and sinusoidal drive and found that such a drive can’t cause a phase transition independent of any frequency regime; it always promotes an area-law entanglement phase in the thermodynamic limit. This result is a bit surprising, especially given that disorder-induced similar-driven systems tend to favor de-localization in the low-frequency regime Sarkar et al. (2021). Our results suggest several interesting directions for future work. The role of disorders, interactions in the monitored Floquet system, and their relationship to Floquet-MBL Sierant et al. (2023) remain a challenging problem.
VI Acknowledgements
R.M. acknowledges the DST-Inspire fellowship by the Department of Science and Technology, Government of India, SERB start-up grant (SRG/2021/002152). The authors also thank M Buchhold for fruitful discussions over private communication on the RG scheme. The authors also thank Xiangyu Cao for the fruitful discussion.
Appendix A $T=100$ result for square pulse protocol
Figure 12: Variation of steady state entanglement entropy with system size for different $\gamma$ values for $T=100$ for square pulse driving protocol with $\epsilon=0$. The dashed lines show the no-drive ($J=1/2$) results.
Here, we discuss the $T=100$ results. Dashed lines correspond to no-drive results for $J=1/2$. It shows that almost up to $L\simeq 100$, the no-drive data is the same as the $T=100$ data. Afterward, $T=100$ data starts bending, while no-drive data keeps on increasing. In the main text, we show that in the case of square pulse protocol and a given driving time period $T$, for $L>>L_{c}$, the steady state entanglement entropy does not change with $L$. The value of $L_{c}$ was estimated using fitting function  (16). It turns out that $L_{c}$ increases exponentially with $T$, the fitted function predict the $L_{c}$ to be ${}10^{16}$ for $T=100$ and $\gamma=0.05$. This suggests that due to our limitations of the accessible system size, we will be unable to identify the area-law phase for $\gamma=0.05$ and $T=100$. As expected, Fig. 12 shows no trace of area law for $\gamma=0.05$. The data can mislead us to predict the usual BKT type critical to area law entanglement phase transition as we increase $\gamma$.
Figure 13: Cost function $F_{Y}$ vs. $\gamma_{c}$ plot for different $\epsilon$ values. The star represents the optimal $\gamma_{c}$ for $\epsilon=5J/4$, and the circle represents the optimal $\gamma_{c}$ value for $\epsilon=3J/4$.
Appendix B Cost-function minimization technique
In order to see the BKT transition, we had to perform a finite-size scaling analysis. The job is to find out the $\gamma_{c}$ for which one can get the best possible data-collapse as observed in Fig. 7 (lower panel). We take $X_{i}=(\gamma-\gamma_{c})(\ln L)^{2}$ for all the values of $\gamma$ and $L$, and $Y_{i}=\overline{S(L,\gamma)}-\overline{S(L,\gamma_{c})}$ sorted according to the increasing values of $X_{i}$. The values of $\overline{S(L,\gamma)}$, which are not in the list, have been calculated using linear extrapolation. Now, one can define the cost function as Šuntajs et al. (2020),
$C_{Y}=\frac{\sum_{i=1}^{P}|Y_{i+1}-Y_{i}|}{\max\{Y_{i}\}-\min\{Y_{i}\}}-1,$
where $P$ is the total number of unique $X_{i}$ present in the list. For a perfect data-collapse, $C_{Y}=0$. We perform a minimization of the function $C_{Y}$ with respect to the free parameter $\gamma_{c}$ and let for $\tilde{\gamma}_{c}$ the cost function is minimum, i.e., $C_{Y}(\tilde{\gamma}_{c})=(C_{Y})_{\text{min}}$ (see Fig. 13). Usually, such minima in $C_{Y}$ may not be that robust, by changing the input data set (adding or removing some data), minima can shift a bit. To incorporate that effect, we identify a set of $\{\gamma_{c}\}$ for which $C_{Y}\leq 2C_{Y}(\tilde{\gamma}_{c})$. Then identify $\gamma^{(1)}_{c}=\text{min}\{\gamma_{c}\}$ (minimum from that set) and $\gamma^{(2)}_{c}=\text{max}\{\gamma_{c}\}$ (maximum from that set). We identify the optimal $\gamma_{c}$ as $\gamma_{c}=(\gamma^{(1)}_{c}+\gamma^{(2)}_{c})/2$ (similar technique was used in Ref. Szyniszewski et al. (2023) to compute the error-bars).
Appendix C F-test
We perform the $F$-test to quantitatively determine the statistical likelihood of our steady state entanglement data presenting a logarithmic scaling or volume law for small $\gamma$. The $F$-test is a likelihood-ratio test that assesses the goodness of fit of two competing statistical models Mood (1950); Xing et al. (2024). For each set of $\overline{S(L)}$ , we produce two curves $\tilde{S}_{L}$ and $\tilde{S}_{\ln(L)}$, which are linear best-fits curves of $\overline{S(L)}$ in $L$ and $\ln(L)$ respectively. These best-fit lines are used to compute the $F$-test statistic,
$F_{T}=\frac{\tilde{E}_{L}}{\tilde{E}_{\ln(L)}},$ (17)
where $\tilde{E}_{L}$ ($\tilde{E}_{\ln(L)}$) is the sum of the squared errors between $\overline{S(L)}$ and $\tilde{S}_{L}$ ($\tilde{S}_{\ln(L)}$). The way $F_{T}$ is constructed, a lower $F_{T}$ will favor volume law, and a higher $F_{T}$ will favor logarithmic scaling. We exercise this test for our $\overline{S(L)}$ vs $L$ data and found the $F_{T}>>1$ for $T=5$, $20$, $100$, $150$, $200$ and for $\gamma=0.05$, and $\epsilon=3J/4$. It confirms that the scaling is most likely logarithmic.